# Ctrl-Room: Controllable Text-to-3D Room Meshes Generation with Layout Constraints

### v1

@@ 3.3 Mask-guided Editing
4 Experiments
We validate our method with experiments and compare it with previous methods on indoor scene generation in accuracy, speed, and result quality. We further show various scene editing results to demonstrate the flexible control of our method.
Table 1: Quantitative Comparison of panorama and mesh generation.
Method Panorama Metrics 2D Rendering Metrics 3D Mesh User Study
FID $\downarrow$
CS $\uparrow$
IS $\uparrow$
Time/s $\downarrow$
CS $\uparrow$
IS $\uparrow$
[…]
3DS $\uparrow$
Text2Light (Chen et al., 2022)
56.22
21.45
4.198
81.56
2.732
2.747
MVDiffusion (Tang et al., 2023c)
34.76
23.93
3.21
208.5
3.27
3.437
Text2Room (Höllein et al., 2023)
$\geq 9,000$
25.90
2.90
2.487
2.588
Ours
21.02
21.58
3.429
61.1 25.97 3.14 3.89 3.746
4.1 Experiment Setup
Dataset: We experiment on the 3D indoor scene dataset Structured3D (Zheng et al., 2020), which consists of 3,500 houses with 21,773 rooms designed by professional artists. Photo-realistic rendered images, such as RGB panorama, semantic panorama, depth map, and normal map are provided in each room. Each room is also labeled with the room layout, represented by the intersection lines of walls, roofs, and floors. 3D bounding boxes of furniture items are provided but noisy. We parse these boxes for all the ‘living rooms’ and ‘bedrooms’. Then, we follow (Wang et al., 2021) to generate text prompts describing the scene layout. Please refer to Appendix Sec.1.3 for more details about data preprocessing. The filtered dataset for training and evaluation includes 4,961 bedrooms and 3,039 living rooms. For each room type, we use $80\%$ of rooms for training and the remaining for testing.
Metrics: We choose Frechet Inception Distance (FID) (Heusel et al., 2017), CLIP Score (CS) (Radford et al., 2021), and Inception Score (IS) (Salimans et al., 2016) to measure the image quality of generated panoramas. We also compare the time cost to synthesize an RGB panorama of size $512\times 1024$. To compare the quality of 3D room models, we follow Text2Room (Höllein et al., 2023) to render images of the 3D room model and measure the CLIP Score (CS) and Inception Score (IS). We further conduct a user study and ask $61$ users to score Perceptual Quality (PQ) and 3D Structure Completeness (3DS) of the final room mesh on scores ranging from 1 to 5.
4.2 Comparison with Previous Methods
Figure 5: Qualitative comparison with previous works. For each method, we show a textured 3D mesh in the first row and a few rendered perspective images in the second row.
Quantitative Comparison To evaluate our generated panoramic images, we follow MVDiffusion (Tang et al., 2023c) to crop perspective images from the 1,181 generated panoramas on the test split and evaluate the FID, CS, IS on the cropped multi-view images. These quantitative results are summarized on the left of Table 1, where our method achieves the best score in FID, which indicates that our method can better capture the room appearance because of its faithful recovery of the room layout. However, our score on CS is slightly lower than MVDiffusion, which seems to be insensitive to the number of objects as we illustrated in the supplementary file, and cannot reflect the quality of room layouts. The IS score depends on the semantic diversity of the cropped images as captured by an image classifier. It turns out Text2Light has the best IS score, since its generated indoor scenes often contain unexpected objects. More panoramic visualization can be found in Appendix Sec.1.5.
In terms of running time, our method takes the shortest time. Averagely speaking, our method takes only 61 seconds to generate a panorama, and another 20 seconds to generate the textured 3D mesh. In comparison, MVDiffusion takes 208 seconds, about 3.5 times longer, to synthesize a panorama. Text2Room needs at least 150 minutes to finish a textured 3D room generation.
We then compare the 3D room models in terms of their rendered images. Because of the expensive running time of Text2Room (Höllein et al., 2023), we only test on 12 examples for this comparison. In this comparison, we further skip Text2light and MVDiffusion since we have compared them on panoramas. As the room layout is better captured with a large FOV, we render 60 perspective images of each scene with a $140^{\circ}$ FOV and evaluate their CS and IS scores respectively. Please refer to Appendix Sec.1.4.1 for more details. The results of this comparison are shown in the middle of Table 1. Our method obtains better scores on both metrics than Text2Room.
We further evaluate the quality of the textured 3D mesh model by user studies. For those panorama generation methods, we utilize the depth estimation work (Shen et al., 2022) to reconstruct a textured 3D room mesh. More details can be found in the Appendix. The results of the user study are shown on the right of Table 1. Users prefer our method over others, for its clear room layout structure and furniture arrangement.
Qualitative Comparison Fig. 5 shows some results generated by different methods. The first row shows a textured 3D room model, and the second row shows some cropped perspective images from the panorama. As we can see, Text2Light (Chen et al., 2022) cannot generate a reasonable 3D indoor scene. It even fails to ensure the loop consistency of the generated panorama, which leads to distorted geometry and room model. Both MVDiffusion (Tang et al., 2023c) and Text2Room Höllein et al. (2023) can generate vivid local images as demonstrated by the perspective renderings in the second row. But they fail to capture the more global scale room layout. Similar effects can be seen from the Fig. 1 (a). These two methods often repeat a dominating object, e.g. a bed in the bedroom or fireplace in the living room, multiple times at different places and violate the room layout constraint. In comparison, our method does not suffer from these problems and generates high-quality results. More examples are provided in the supplementary file.
Figure 6: Room editing examples. (a) resize the TV, (b) replace the chair with a new one.
4.3 Interactive Scene Editing
We demonstrate the scene editing capability of our method in  Fig. 6. In this case, we resize the TV and replace the chair in the generated results. Fig. 1 (b) shows two additional examples of replacing the TV and TV stand and moving the picture upwards. Our method can keep the visual appearance of the moved/resized objects unchanged after editing. More examples can be found in the supplementary file.
5 conclusion and discussion
We present Ctrl-Room, a flexible method to achieve editable and structurally plausible 3D indoor scene generation. It consists of two stages, the layout generation stage and the appearance generation stage. In the layout generation stage, we design a scene code to parameterize the scene layout and learn a text-conditioned diffusion model for text-driven layout generation. In the appearance generation stage, we fine-tune a ControlNet model to generate a vivid panorama image of the room with the guidance of the layout. Finally, a high-quality 3D room with a structurally plausible layout and realistic textures can be generated. We conduct extensive experiments to demonstrate that Ctrl-Room outperforms existing methods for 3D indoor scene generation both qualitatively and quantitatively, and supports interactive 3D scene editing.

### v3

@@ 9 Mask-Guided Editing
10 Dataset
Structured3D dataset preprocessing Structured3D consists of $3,500$ houses with $21,773$ rooms, where each room is designed by professional designers with rich 3D structure annotations, including the room planes, lines, junctions, and orientated bounding box of most furniture, and photo-realistic 2D renderings of the room. In our work, we use the 3D orientated bounding boxes of furniture, 2D RGB panorama, and 3D lines and planes of each room. While the original dataset lacks semantic class labels for each furniture bounding box. The dataset preprocessing aims to produce clean ground truth data for our layout generation module and appearance generation module.
  • Orientated Object Bounding Box Annotation. As the original dataset lacks semantic label for each orientated object bounding box, we first unproject the RGB panorama and depth map into a point cloud of the room, then manually annotate the object semantic class and add more accurate object bounding boxes based on the noisy annotation of the original version. As shown in Fig. 12, by using labelCloud [26], three data annotators worked for 1200 hours to annotate 5,064 bedrooms, 3,064 livingrooms, 2,289 kitchens, 698 studies, and 1,500 bathrooms, getting nearly 150K accurate orientated 3D bounding boxes across 25 object categories.
  • Scene Node Encoding. We define our holistic scene code based on a unified encoding of walls and object bounding box. Each object $o_{j}$ is treated as a node with various attributes, i.e., center location $l_{i}\in\mathbb{R}^{3}$, size $s_{i}\in\mathbb{R}^{3}$, orientation $r_{i}\in\mathbb{R}$, class label $c_{i}\in\mathbb{R}^{C}$. The orientated bounding box is off-the-shelf, we extract the inner walls based on the line junctions and corners of the 3D room. Then we put the orientated object bounding boxes and walls into a compact scene code. Concretely, we define an additional ’empty’ object and pad it into scenes to have a fixed number of object across scenes. Each object rotation angle is parametrized by a 2-d vector of cosine and sine values. Finally, each node is characterized by the concatenation of these attributes as $\mathbf{o}_{i}=[c_{i},l_{i},s_{i},{\rm cos}r_{i},{\rm sin}r_{i}]$.
  • data filtering. We start by filtering out those problematic scenes such as rooms with wall number less than $4$ or larger than $20$. We also remove those scenes with too few or too many objects. The number of walls of valid bedrooms is between $4$ and $10$, and that of objects is between $3$ and $13$. As for living rooms, the minimum and maximum numbers of walls are set to $4$ and $20$, and that of objects are set to $3$ and $24$ respectively. The number of walls for valid kitchens, studies, and bathrooms is the same as for bedrooms, while the objects number is between $3$ and $24$. Thus, the number of scene nodes is $N=23$ in bedrooms, $N=44$ in living rooms, and $N=34$ in kitchens, studies, and bathrooms. After filtering, we get 4,961 bedrooms, 3,039 living rooms, 1,848 kitchens, 638 studies, and 1,356 bathrooms.
Figure 12: Example of object bounding box annotation.
Text Prompt Generation We follow the SceneFormer [42] to generate text prompts describing partial scene configurations. Each text prompt contains two to four sentences. The first sentence describes how many walls are in the room, then the second sentence describes two or three existing furniture in the room. The following sentences mainly describe the spatial relations among the furniture, please refer to SceneFormer [42] and DiffuScene [37] for more detailed explanation of relation-describing sentences. In this way, we can get some relation-describing sentences to depict the partial scene. Finally, we randomly sampled zero to two relation-describing sentences to form the text prompt for 3D room generation.
11 Implementation details
Training and inference details.
  • In the layout generation stage, We train the scene code diffusion model on our processed typical indoor rooms data of Structured3D [49] for $200,000$ steps. The frozen text encoder we adopted is the same as Stable Diffusion [25]. The training is performed using the AdamW optimizer with a batch size of $128$ and a learning rate of $41e-$ 4, utilizing $2$ A6000 GPUs. During the inference process, we utilize the DDIM [34] sampler with a step size of $200$ to perform scene code denoising.
  • In the appearance generation stage, we fine-tune the segmentation-conditional ControlNet model based on the pairwise semantic and RGB panorama of Structured3D. The fine-tuning process is implemented on two A6000 GPUs for $150$ epochs(about 3 days). In the inference phase, we generate high-fidelity and loop-consistent RGB panorama through DDIM sampler with $100$ steps, rotating both semantic layout panorama and the denoised image for $\mathbf{\gamma}=90^{\circ}$ at each step.
  • As for the mask-guided editing module, we utilize the fine-tuned Control-Seg model to inpaint the background content and optimize the latents of the edited panorama. In inpainting step, the weights used too fuse the unpainted area and unchanged area are set $\lambda_{\rm ori}=0.8,\lambda_{\rm new}=0.2$ In the optimization step, the maximum iteration is $M=50$, the learning rate $\eta$ for optimization is initialized to $0.1$ and then gradually decreases to $0.01$.
Table 3: The computational cost comparison on A6000. Since Text2Room does not offer a standardized neural network, we cannot measure its parameters.
Method Inference Time/s GPU Memory Params
Text2Light 81.56 5.46G 630.66M
MVDiffusion 208.5 8.74G 1352.54M
Text2Room $>9,000$ $>16$G -
Ours 61.1 1.95G + 10.41G 63.51M + 1220.62M
11.1 Baseline Implementations
We provide implementation details for baseline methods in the following:
  • MVDiffusion [39]: To get a high-resolution photo realistic panorama, MVDiffusion employs $8$ branches of SD [25] model and correspondence-aware attention mechanism to generate multi-view images simultaneously. We first fine-tune the pre-trained model of MVDiffusion on Structured3D for 10 epochs(about 3 days). Since each generated subview image of MVDiffusion is at $512\times 512$ resolution, the final panorama is pretty large. We resize the generated panorama of MVDiffusion from $4096\times 2048$ to $1024\times 512$. Then the 8 subview perspective images are extracted from the post-processed panorama using the same camera settings (FOV=$90^{\circ}$,rotation=$45^{\circ}$). The same operation is adopted on our generated panoramic images. Finally, we combine the panorama from MVDiffusion with our panoramic reconstruction module to create a 3D mesh.
  • Text2Light [6]: Text2Light creates HDR panoramic images from text using a multi-stage auto-regressive generative model. We choose Text2Light as one of the baseline for our panorama generation and 3D room mesh generation. We first generate RGB panoramas from the input text using Text2Light, then lift it into 3D mesh using the same panoramic reconstruction module as our method. When evaluate the panoramic image quality, we adopt the same processing as MVDiffusion to get multi-view perspective images of Text2Light.
  • Text2Room [12]: Text2Room is the current state-of-the-art and off-the-shelf method for 3D room mesh generation. It utilizes 20 camera spots of a pre-defined trajectory to expand new areas as much as possible by generating 10 images at each spot. Here We use its final fused poison mesh for 3D mesh comparison. For a fair comparison of 2D renderings evaluation, we only use the renderings at the origin of the final mesh.
  • Text2NeRF [47] generates 3D scenes from a text prompt using NeRF as the 3D representation and leverages a pre-trained text-to-image diffusion model and monocular depth estimation to constrain the 3D reconstruction. However, we found it fails to reconstruct $360^{\circ}$ scenes. We present some NeRF reconstructions from Text2NeRF stitched into panorama images in Fig.19. Note that only $\sim$154° horizontal field of view (FOV) and $\sim$113° vertical FOV is shown since the rest of the scene is not reconstructed by the method. Thus we skip the comparison with this method.
We also compare the model complexity between ours and the baselines, Table 3 shows our method achieves significantly faster runtime for generating a 3D room compared to existing methods. The difference in GPU memory footprint and parameter quantity is not significant. Despite having two diffusion models, our layout generation stage only needs to produce high-level room layouts and furniture arrangements, resulting in low computational costs and model complexity.
12 Panorama Generation Comparison
Fig. 13 presents additional results for panorama generation. Given a simple partial-scene text prompt, our approach obtains better RGB panorama than that of Text2Light [6] and MVDiffusion [39], which demonstrates the effectiveness of our well-designed framework. While Text2Light suffers from the inconsistent loop and unexpected content of the generated panorama, MVDiffusion fails to recover a reasonable room layout from the text prompt.
Figure 13: Qualitative comparison for panorama generation. Generated panorama is visualized in a panoramic image viewer to facilitate the user to check the global content of panorama. The left side of each column is two zoom-in views, and the right side is the fisheye view. Text2Light [6] exists serious inconsistent problem on the border of the generated panorama, it also shows a lot of unexpected stuff in the image. MVDiffusion [39] fails to synthesize reasonable content for the target room type. In contrast, our method obtains layout plausible and vivid panorama from the given text prompt of partial scene.
13 Additional Qualitative Results
In Fig. 14, we first visualize more generated room layouts generation of typical rooms in the format of semantic 3D bounding boxes. Then, we show additional qualitative comparison results between our method and baselines in Fig. 16. We demonstrate more scene editing results of our method in Fig. 15.
Figure 14: Additional room layout generations. In the bedroom, the bed is often attached to the wall, with a picture above it and a television in front of it. In the living room, there is often a double-seat sofa accompanied by a table and a single-seat sofa. The dining table is usually placed in a separate area of the living room, along with cabinets and chairs. In the kitchen, common furniture includes a stove, sink, fridge, and hood, which are all well-placed in the room. In the study, there is typically a desk accompanied by a chair and one or more bookshelves, and sometimes there is also a bed in the room. In the bathroom, there is usually a sink with a mirror, a toilet, and a shower.
Figure 15: Additional scene editing results. In each sub-figure, the left part is the original 3D room, the right part shows the final mesh after users’ interactive editing.
Figure 16: Additional qualitative comparison with previous works.
14 User Study
Follow Text2Room [12], we conduct a user study and ask $n=61$ ordinary users to score the Perceptual Quality(PQ) and 3D Structure Completeness(3DS) of the generated room on a scale of $1-5$. Different from Text2room which only demonstrates the perspective renderings of the 3D room, we directly show users the generated mesh to get a global evaluation of the whole generated 3D room. We show an example of the presented interface of the user study in Fig. 17. In total, we presented $40$ top-down views from $10$ scenes and report averaged results for each method. Users favor our approach, which emphasizes the superiority of our more plausible geometry, along with the vivid texture.
Figure 17: User study interface. We provide users with multiple top-down images from different methods and ask users to rate the given 3D meshes on a scale from 1 to 5, according to the criteria of Perceptual Quality and 3D Structure Completeness.
15 Free style prompts
We show the adaptability of our method by utilizing Large Language Model (LLM) GPT-4 Vision (GPT-4V) [46] to generate text captions from panorama images of Structured3D [49] bedroom scenes. The prompt used for the LLM is as shown in Table 4.
We train and test with the LLM generated captions as conditioning for layout generation. Fig. 18 shows some results from the test set and corroborates our ability to produce plausible 3D room layout following free-style test prompts.
Figure 18: Text-conditioned layout generation on Structured3D using GPT-4V text prompts. Our method synthesizes a plausible scene layout that matches the description.
Table 4: Prompt for GPT-4V to generate captions from panorama images
Describe what is displayed in the panoramic image succinctly in 3 or 4 sentences encoded in ASCII. Do not use lengthy or compound sentences. Do not mention that it is an image or a panoramic image. Do not describe the background, lighting, color palette or count the number of objects. Do not describe size like “small”, “large”, etc. Describe the relative positions of each objects in the scene using only these relationships: “on”, “above”, “surrounding”, “inside”, “left touching”, “right of”, “front touching”, “in front of”, “right touching”, “left of”, “behind touching”, “behind”, “next to”, “left of”, “right of”. Optionally, describe the object attributes (color, texture etc). In the description only use these objects: table, night stand, picture, door, cabinet, curtain, bathtub, bed, sink, fridge, shelves, window, lamp, chair, pillow, dresser, bookshelf, sofa, counter, desk, mirror, television, wall
Figure 19: Text2NeRF results. The NeRF reconstructions are stitched into panorama images. Only 154° horizontal FOV and 113° vertical FOV is shown since the method was not able to reconstruct the rest of the scene.